A solution of Na3AsO4 is added dropwise to a solution that is 0.0220 M in Cu2+ and 0.492 M in Ag+.

The Ksp of Cu3(AsO4)2 is 7.95e-36.
The Ksp of Ag3AsO4 is 1.03e-22.






(a) What concentration of AsO43- is necessary to begin precipitation? (Neglect volume changes.)

[AsO43-] =  M.



(b) Which cation precipitates first?

Cu2+

Ag+    






(c) What is the concentration of AsO43- when the second cation begins to precipitate?

[AsO43-] =  M.
Answer #1

a) Solubility equilibrium of Cu3(AsO4)2 is

Cu3(AsO4)2(s) <-------> 3Cu2+(aq) + 2AsO43-(aq)

Ksp = [Cu2+]3[AsO43-]2 = 7.95 ×10-36

substituting concentration of Cu2+ in Ksp expression

(0.0220M)3 × [AsO43-]2 = 7.95 ×10-36M5

[AsO43-]2 = 7.47 × 10-31M2

[AsO43-] = 8.64 ×10-16M

concentration of AsO43- required to start precipitation of Cu3(AsO4)2 = 8.64 ×10-16M

Solubility equilibrium of Ag3AsO4 is

Ag3AsO4(s) <--------> 3Ag+(aq) + AsO43-(aq)

Ksp = [Ag+]3[AsO​​​​​​43-] = 1.03 ×10-22

substitute the concentration of Ag+ in Ksp expression

( 0.492M)3 × AsO43- = 1.03 ×10-22M4

AsO43- = 8.65 × 10-22M

Therefore,

Concentration of AsO43- necessary to begin precipitation = 8.65 ×10-22M

b)

Concentration of AsO43- necessary to start precipitation of Ag3AsO4 is lower than concentration of AsO43- necessary to start precipitation of Cu3(AsO4)2

Therefore,

Ag+ precipitate first

c)

Substitute the concentration of AsO43 necessary to start precipitation of Cu3(AsO4)2in Ksp expression of Ag3AsO4

[Ag+]3 × [AsO43-] = 1.03 ×10-22

[Ag+]3 × ( 8.64 × 10-16M ) = 1.03 ×10-22M4

[Ag+]3 = 1.192× 10-7

[Ag+] = 4.92 × 10-3M

Therefore,

concentration of AsO43- when second cation begins to precipitate = 4.92 × 10-3M